### v1

# Mini-batch Submodular Maximization

Abstract
We present the first mini-batch algorithm for maximizing a non-negative monotone decomposable submodular function, $F=\sum_{i=1}^{N}f^{i}$, under a set of constraints. We improve over the sparsifier based approach both in theory and in practice. We experimentally observe that our algorithm generates solutions that are far superior to those generated by the sparsifier based approach.
1 Introduction
Submodular functions capture the natural property of diminishing returns which often arises in machine learning, graph theory and economics. For example, imagine you are given a large set of images, and your goal is to extract a small subset of images, that best represent the original set (e.g., creating thumbnails for a Youtube video). Intuitively, this is a submodular optimization problem, as the more thumbnails we have, the less we gain by adding an additional thumbnail.
Formally, a set function $F:2^{E}\rightarrow\mathbb{R}^{+}$, on a ground set $E$, is submodular if for any subsets $S\subseteq T\subseteq E$ and $e\in E\setminus T$, it holds that
$F(S+e)-F(S)\geq F(T+e)-F(T)$
Decomposable submodular functions
In many natural scenarios $F$ is decomposable: $F=\sum_{i=1}^{N}f^{i}$, where each $f^{i}:2^{E}\rightarrow\mathbb{R}^{+}$ is a non-negative submodular function on the ground set $E$ with $|E|=n$. That is, $F$ can be written as a sum of “simple” submodular functions. We assume that every $f^{i}$ is represented by an evaluation oracle that when queried with $S\subseteq E$ returns the value $f^{i}(S)$. Our goal is to maximize $F$ under some set of constraints while minimizing the number of oracle calls to $\left\{f^{i}\right\}$. An excellent survey of the importance of decomposable functions is given in Rafiey and Yoshida (2022), which we summarize below.
Decomposable submodular functions are prevalent in both machine learning and economic studies. In economics, they play a pivotal role in welfare optimization during combinatorial auctions (Dobzinski and Schapira, 2006; Feige, 2009; Feige and Vondrák, 2006; Papadimitriou et al., 2008; Vondrák, 2008). In machine learning, these functions are instrumental in tasks like data summarization, aiming to select a concise yet representative subset of elements. Their utility spans various domains, from exemplar-based clustering by (Dueck and Frey, 2007) to image summarization (Tschiatschek et al., 2014), recommender systems (Parambath et al., 2016) and document summarization (Lin and Bilmes, 2011). The optimization of these functions, especially under specific constraints (e.g., cardinality, matroid) has been studied in various data summarization settings (Mirzasoleiman et al., 2016a, b, c) and differential privacy (Chaturvedi et al., 2021; Mitrovic et al., 2017; Rafiey and Yoshida, 2020).
In many of the above applications $N$ (the number of underlying submodular functions) is extremely large, making the evaluation of $F$ prohibitively slow. We illustrate this with a simple example.
Example (Welfare maximization)
Imagine you are tasked with deciding on a meal menu for a large group of $N$ people (e.g., all students in a university, all high school students in a country). You need to choose $k$ ingredients to use from a predetermined set (chicken, fish, beef, etc.) of size $n$. Every student has a specific preference, modeled as a monotone submodular function $f^{i}$. Our goal is to maximize the “social welfare”, $F(S)=\sum f^{i}(S)$, over all students.
The greedy algorithm
Going forward we focus on the case where $\forall i,f^{i}$ is monotone ($\forall S\subseteq T\subseteq E,f^{i}(T)\geq f^{i}(S)$), which in turn means $F$ is monotone. This is applicable to the above example (e.g., students are happier with more food varieties).
For ease of presentation let us first focus on maximizing $F$ under a cardinality constraint $k$, i.e., $\max F(S),\left|S\right|\leq k$. The classical greedy algorithm (Nemhauser et al., 1978) (Algorithm 1) achieves an optimal $(1-1/e)$-approximation for this problem. For $S,A\subseteq E$ we define $F_{S}(A)=F(S+A)-F(S)$. We slightly abuse notation and write $F_{S}(e),F(e)$ instead of $F_{S}(\left\{e\right\}),F(\left\{e\right\})$.
1 $S_{1}\leftarrow\emptyset$
2 for $j=1$ to $k$ do
3       $e^{\prime}=\arg\max_{e\in E\setminus S_{j}}F_{S_{j}}(e)$
4       $S_{j+1}=S_{j}+e^{\prime}$
5
6 end for
7return $S_{k+1}$
Algorithm 1 Greedy submodular maximization under a cardinality constraint
When $F$ is decomposable, and each evaluation of $f^{i}$ is counted as an oracle call, the above algorithm requires $O(Nnk)$ oracle calls. This can be prohibitively expensive if $N\gg n$. Looking back at our example, Algorithm 1 will require asking every student $k$ questions. That is, we would start by asking all students: “Rate how much you would like to see food $x$ on the menu” for $n$ possible options. We would then need to wait for all of their replies, and continue with “Given that chicken is on the menu, rate how much you would like to see food $y$ on the menu” for $n-1$ possible options and so on. Even if we do an online poll, this is still very time-consuming (as we must wait for everyone to reply in each of the $k$ steps).
If we could sample a representative subset of the students, we could greatly speed up the above process. Specifically, we would like to eliminate the dependence on $N$ in Algorithm 1.
A sparsifier based approach
Recently, Rafiey and Yoshida (2022) showed how to construct a sparsifier for $F$. That is, given a parameter ${\epsilon}>0$ they show how to find a vector $w\in\mathbb{R}^{N}$ such that the number of non-zero elements in $w$ is small in expectation and the function $\hat{F}=\sum_{i=1}^{N}w_{i}f^{i}$ satisfies with high probability (w.h.p) [^1] Probability at least $1-1/n^{c}$ for an arbitrary constant $c>1$. The value of $c$ does affect the asymptotics of the results we state (including our own). that
$\forall S\subseteq E,(1-{\epsilon})F(S)\leq\hat{F}(S)\leq(1+{\epsilon})F(S)$
Specifically, every $f^{i}$ is sampled with probability $\alpha_{i}$ proportional to $p_{i}=\max_{S\subseteq E,F(S)\neq 0}\frac{f^{i}(S)}{F(S)}$. If it is sampled, it is included in the sparsifier with weight $1/\alpha_{i}$, which implies that $\mathbb{E}\left[{w_{i}}\right]=1$. While calculating the $p_{i}$’s exactly requires exponential time, Rafiey and Yoshida (2022) make do with an approximation, which can be calculated using interior point methods (Bai et al., 2016).
Rafiey and Yoshida (2022) show that if all $f^{i}$’s are non-negative and monotone [^2] Rafiey and Yoshida (2022) also present results for non-monotone functions, however, Kudla and Zivný (2023) point out an error in their calculation and note that the results only hold when all $f^{i}$’s are monotone., the above sparsifier can be constructed by an algorithm that requires $poly(N)$ oracle evaluations and the sparsifier will have expected size $O({\epsilon}^{-2}Bn^{2.5}\log n)$, where $B=\max_{i\in[N]}B_{i}$ and $B_{i}$ is the number of extreme points in the base polyhedron of $f^{i}$. They extend their results to matroid constraints of rank $r$ and show that a sparsifier with expected size $O({\epsilon}^{-2}Brn^{1.5}\log n)$ can be constructed.
For the specific case of a cardinality constraint $k$, this implies a sparsifier of expected size $O({\epsilon}^{-2}Bkn^{1.5}\log n)$ can be constructed using $poly(N)$ oracle evaluations. The sparsifier construction is treated as a preprocessing step (we elaborate on this in Section 1.1), and therefore the actual execution of Algorithm 1 on the sparsifier requires only $O({\epsilon}^{-2}Bk^{2}n^{2.5}\log n)$ oracle evaluations to get a $(1-1/e-{\epsilon})$ approximation. This is an improvement over Algorithm 1 when $N\gg n,B$.
Recently, Kudla and Zivný (2023) showed improved results for the case of bounded curvature. The curvature of a submodular function $F$ is defined as $c=1-\min_{S\subseteq E,e\in E\setminus S}\frac{F_{S}(e)}{F_{\emptyset}(e)}$. We say that $F$ has bounded-curvature if $c<1$. Submodular functions with bounded curvature (Conforti and Cornuéjols, 1984) offer a balance between modularity and submodularity, capturing the essence of diminishing returns without being too extreme.
They show that when the curvature of all $f^{i}$’s and of $F$ is constant it is possible to reduce the preprocessing time to $O(Nn)$ oracle queries and to reduce the size of the sparsifier by a factor of $\sqrt{n}$. Furthermore, their results extend to the much more general case of $k$-submodular functions. While this significantly improves over the number of oracle calls compared to (Rafiey and Yoshida, 2022), the running time of the preprocessing step depends on $\log\left(\max_{i\in[N]}\frac{\max_{e\in E}f^{i}_{\emptyset}(e)}{\min_{e\in E,f^{i}_{\emptyset}(e)>0}f^{i}_{\emptyset}(e)}\right)$.
The current state of the art is due to Kenneth and Krauthgamer (2023) where they show that by sampling according to $p_{i}=\max_{e\in E,F(e)\neq 0}\frac{f^{i}(e)}{F(e)}$ it is possible to get both a fast sparsifier construction time of $O(Nn)$ and a small sparsifier size of $O(\frac{n^{3}}{{\epsilon}^{2}})$. Their analysis also implies that if the solution size is bounded by $k$ (e.g., a cardinality constraint) a sparsifier of size $O(\frac{nk^{2}}{{\epsilon}^{2}})$ is sufficient. They also present results for general submodular functions, however we only focus on their results for monotone functions which are relevant for this paper.
1.1 Our results
In this work, we focus on the greedy algorithm for constrained submodular maximization. We show that instead of sparsifying $F$, better results can be achieved by using mini-batches during the execution of the greedy algorithm. That is, rather than sampling a large sparsifier $\hat{F}$ and performing the optimization process, we show that if we sample a much smaller sparsifier (a mini-batch), $\hat{F}^{j}$, for the $j$-th step of the greedy algorithm, we can achieve improved results both in theory (Table 1) and in practice (Section 3). Most notably, we observe that the mini-batch approach is superior to the sparsifier approach for small batch values on various real world datasets.
While the mini-batch approach results in an improvement in performance, the sparsifier approach has the benefit of being independent of the algorithm. That is, while any approximation algorithm executed on a sparsifier immediately achieves (nearly) the same guarantees for the original function, we need to re-establish the approximation ratio of our mini-batch algorithm for different constraints. Although these proofs are often straightforward, compiling an exhaustive list of where the mini-batch method is applicable is both laborious and offers limited insights.
To illustrate the effectiveness of our method while maintaining readability, we focus on two widely researched constraints: the cardinality constraint and the $p$-system constraint (defined later in the section). The cardinality constraint was chosen for its simplicity and its prominence in research, while the $p$-system constraint was chosen for its broad applicability. We strongly believe that our approach could be applied beyond submodular functions (e.g., $k$-submodular functions, similar to Kudla and Zivný (2023)), achieve better approximation guarantees for specific constraints, and even applied beyond the greedy algorithm.
We compare our results with the results of Rafiey and Yoshida (2022); Kudla and Zivný (2023); Kenneth and Krauthgamer (2023) and the naive algorithm (without sampling or sparsification) in Table 1333Where $\widetilde{O}$ hides $\log n$ factors.. While our results hold for the unbounded curvature case, we can get improved performance if the curvature is bounded. It’s worth noting that while Kudla and Zivný (2023) assume every $f^{i}$ has bounded curvature, we only require $F$ to have bounded curvature.
Preprocessing Oracle queries
Unbounded Curvature
Naive None $O(Nnk)$
Rafiey and Yoshida $poly(N)$ $\widetilde{O}(B\cdot\frac{k^{2}n^{2.5}}{{\epsilon}^{2}})$
Kenneth and Krauthgamer $O(Nn)$ $\widetilde{O}(\frac{k^{3}n^{2}}{{\epsilon}^{2}})$
Our results $O(Nn)$ Cardinality: $\widetilde{O}\left(\frac{k^{2}n^{2}}{{\epsilon}^{2}}\right)$ $p$-system: $\widetilde{O}\left(\frac{k^{2}pn^{2}}{{\epsilon}^{2}}\right)$
Bounded Curvature
Kudla and Zivný $O(Nn)$ $\widetilde{O}(B\cdot\frac{k^{2}n^{2}}{{\epsilon}^{2}})$
Our results $O(Nn)$ $\widetilde{O}(\frac{kn^{2}}{{\epsilon}^{2}})$
Table 1: Comparison of the number of oracle queries during preprocessing and during execution. Results are for the greedy algorithm under both a cardinality constraint ($(1-1/e-\epsilon)$-approximation) and a $p$-system constraint ($(\frac{1-{\epsilon}}{p+1})$-approximation). Unless explicitly stated the number of queries is the same for both constraints. For ease of presentation and to allow comparison with Kudla and Zivný, we assume the curvature is constant. The preprocessing step of Rafiey and Yoshida uses interior point methods, therefore, it is significantly more costly than $O(Nn)$.
Meta greedy algorithm
Our starting point is the meta greedy algorithm (Algorithm 2). The algorithm executes for $k\leq n$ iterations where $k$ is some upper bound on the size of the solution. At every iteration, the set $A_{j}\subseteq E\setminus S_{j}$ represents some constraint that limits the choice of potential elements to extend $S_{j}$. The algorithm terminates either when the solution size reaches $k$ or when no further extensions to the current solution are possible (i.e., $A_{j}=\emptyset$). Furthermore, the algorithm does not have access to the exact incremental oracle, $F_{S_{j}}$, at every iteration, but only to some approximation (which may differ between iterations).
1 $S_{1}\leftarrow\emptyset$
2 Let $k$ be an upper bound on the size of the solution
3 for $j=1$ to $k$ do
4       Let $A_{j}\subseteq E\setminus S_{j}$              $\triangleright$ Problem specific constraint (e.g., $A_{j}=E\setminus S_{j}$ for card. constraint)
5       if $A_{j}=\emptyset$ then return $S_{j}$ Let $\hat{F}^{j}_{S_{j}}$ be an approximation for $F_{S_{j}}$    $\triangleright$ Problem specific approximation
6       $e_{j}=\arg\max_{e\in A_{j}}\hat{F}^{j}_{S_{j}}(e)$
7       $S_{j+1}=S_{j}+e_{j}$
8
9 end for
return $S_{k+1}$
Algorithm 2 Meta greedy algorithm with an approximate oracle
Before we formally define “approximation” in the above, let us note that when we have access to exact values of $F_{S_{j}}$, Algorithm 2 captures many variants of the greedy submodular maximization algorithm. For example, setting $A_{j}=E\setminus S_{j}$ we get the algorithm of Nemhauser et al. (1978) for maximizing a non-negative submodular function under a cardinality constraint. This meta-algorithm also captures the case of maximization under a $p$-system constraint.
$p$-systems
The concept of $p$-systems offers a generalized framework for understanding independence families, parameterized by an integer $p$. We can define a $p$-system in the context of an independence family $\mathcal{I}\subseteq 2^{E}$ and $E^{\prime}\subseteq E$. Let $\mathcal{B}(E^{\prime})$ be the maximal independent sets within $\mathcal{I}$ that are also subsets of $E^{\prime}$. Formally,
$\mathcal{B}(E^{\prime})=\{A\in\mathcal{I}|A\subseteq E^{\prime}\text{ and no }A^{\prime}\in\mathcal{I}\text{ exists such that }A\subset A^{\prime}\subseteq E^{\prime}\}.$
A distinguishing characteristic of a $p$-system is that for every $E^{\prime}\subseteq E$, the ratio of the sizes of the largest to the smallest sets in $\mathcal{B}(E^{\prime})$ does not exceed $p$:
$\frac{\max_{A\in\mathcal{B}(E^{\prime})}|A|}{\min_{A\in\mathcal{B}(E^{\prime})}|A|}\leq p$
The significance of $p$-systems lies in their ability to encapsulate a variety of combinatorial structures. For instance, when we consider the intersection of $p$ matroids, they can be aptly described using $p$-systems. To provide more tangible examples, in graph theory, the collection of matchings in a standard graph can be viewed as a 2-system. Extending this to hypergraphs, where edges might have cardinalities up to $p$, the set of matchings therein can be conceptualized as a $p$-system.
The greedy algorithm for $p$-systems
Formally, the optimization problem can be expressed as: $\max_{S\in\mathcal{I}}F(S)$ where the pair $(E,\mathcal{I})$ characterizes a $p$-system and $F:2^{E}\rightarrow\mathbb{R}^{+}$ denotes a non-negative monotone submodular set function. It was shown by Nemhauser et al. (1978) that the natural greedy approach achieves an optimal approximation ratio of $\frac{1}{p+1}$. Setting $A_{j}=\left\{e\mid S_{j}+e\in\mathcal{I}\right\}$ (i.e., $S_{j}$ remains an independent set after adding $e$) in Algorithm 2 we get the greedy algorithm of Nemhauser et al. (1978). Note that for general $p$-systems it might be that $k=n$, however, there are very natural problems where $k\ll n$. For example, for maximum matching $E$ corresponds to all edges in the graph, which can be quadratic in the number of nodes, while the solution is at most linear in the number of nodes.
Approximate oracles
In many scenarios we do not have access to exact values of $F_{S_{j}}$, and instead we must make do with an approximation. We start with the notion of an approximate incremental oracle introduced in (Goundan and Schulz, 2007). We say that $\hat{F}^{j}_{S_{j}}$ is an $(1-{\epsilon})$-approximate incremental oracle if
$\forall e\in A_{j},(1-\epsilon)F_{S_{j}}(e)\leq\hat{F}^{j}_{S_{j}}(e)\leq(1+\epsilon)F_{S_{j}}(e)$
It was shown in (Goundan and Schulz, 2007; Călinescu et al., 2011) [^4] Strictly speaking, both Goundan and Schulz (2007) and Călinescu et al. (2011) define the approximate incremental oracle to be a function that returns $e_{j}$ at iteration $j$ of the greedy algorithm such that $\forall e\in A_{j},F_{S_{j}}(e_{j})\geq(1-{\epsilon})F_{S_{j}}(e)$. Our definition guarantees this property while allowing easy analysis of the mini-batch algorithm. that given a $(1-{\epsilon})$-approximate incremental oracle, the greedy algorithm under both a cardinality constraint and a $p$-system constraint achieves almost the same (optimal) approximation ratio as the non-approximate case.
Theorem 1.
Algorithm 2 with an $(1-\epsilon)$-approximate incremental oracle has the following guarantees w.h.p.
  • It achieves a $(1-1/e-{\epsilon})$-approximation under a cardinality constraint $k$ (Goundan and Schulz, 2007).
  • It achieves a $(\frac{1-{\epsilon}}{1+p})$-approximation under a $p$-system constraint (Călinescu et al., 2011).
We introduce a weaker type of approximate incremental oracle, which we call an additive approximate incremental oracle. We extend the results of Theorem 1 for this case. Let $S^{*}$ be some optimal solution for $F$ (under the relevant set of constraints). We say that $\hat{F}^{j}_{S_{j}}$ is an additive ${\epsilon}^{\prime}$-approximate incremental oracle if
$\forall e\in A_{j},F_{S_{j}}(e)-\epsilon^{\prime}F(S^{*})\leq\hat{F}^{j}_{S_{j}}(e)\leq F_{S_{j}}(e)+\epsilon^{\prime}F(S^{*})$
This might seem problematic at first glance, as it might be the case that $F(S^{*})\gg F_{S_{j}}(e)$. Luckily, the proofs guaranteeing the approximation ratio are linear in nature. Therefore, by the end of the proof we end up with an expression of the form:
$F(S_{k+1})\geq F(S^{*})\beta-\gamma{\epsilon}^{\prime}F(S^{*})$
Where $\beta$ is the desired approximation ratio and $\gamma$ depends on the parameters of the problem (e.g., $\beta=(1-1/e),\gamma=2k$ for a cardinality constraint). We can achieve the desired result by setting ${\epsilon}^{\prime}={\epsilon}/\gamma$. We state the following theorem (the proofs are very similar to those of Goundan and Schulz (2007); Călinescu et al. (2011), and we defer them to the Appendix).
Theorem 2.
Algorithm 2 with an additive ϵsuperscriptitalic-${\epsilon}^{\prime}$-approximate incremental oracle has the following guarantees w.h.p.
  • If ${\epsilon}^{\prime}<{\epsilon}/2k$, it achieves a $(1-1/e-{\epsilon})$-approximation under a cardinality constraint $k$.
  • If ${\epsilon}^{\prime}<{\epsilon}/2kp$, it achieves a $(\frac{1-{\epsilon}}{1+p})$-approximation under a $p$-system constraint.
Mini-batch sampling
Our main result shows that when $\hat{F}^{j}_{S_{j}}$ is sampled using mini-batch sampling we indeed get, w.h.p, an (additive) approximate incremental oracle for every step of the algorithm. We present our sampling procedure in Algorithm 3 and the complete mini-batch algorithm in Algorithm 4.
Similar to Rafiey and Yoshida (2022); Kenneth and Krauthgamer (2023) we treat the computation of the $p_{i}$’s as a preprocessing step. The justification for this, is that the $p_{i}$’s do not depend on the constraints of the problem. Therefore, computing the $p_{i}$’s a single time, we can execute our algorithm on various constraints (e.g., different $p$-systems).
1 $w\leftarrow\vec{0}$ for  $i=1$ to $N$  do
2       $\alpha_{i}\leftarrow\min\{1,\alpha p_{i}\}$ $w_{i}\leftarrow 1/\alpha_{i}$ with probability $\alpha_{i}$                     $\triangleright$ Do nothing with probability $1-\alpha_{i}$
3 end for
4return $\hat{F}=\sum_{i=1}^{N}w_{i}f^{i}$
Algorithm 3 Sample($\alpha,\left\{p_{i}\right\}_{i=1}^{N}$)
1 $\forall i\in[N],p_{i}\leftarrow\max_{e\in E,F_{\emptyset}(e)\neq 0}\frac{f^{i}_{\emptyset}(e)}{F_{\emptyset}(e)}$                    $\triangleright$ Preprocessing
2 $\alpha$ is a batch parameter
3 $S_{1}\leftarrow\emptyset$
4 Let $k$ be an upper bound on the size of the solution
5 for $j=1$ to $k$ do
6       Let $A_{j}\subseteq E\setminus S_{j}$                                         $\triangleright$ Problem specific constraint
7       if $A_{j}=\emptyset$ then return $S_{j}$ Let $\hat{F}^{j}_{S_{j}}\leftarrow Sample(\alpha,\left\{p_{i}\right\}_{i=1}^{N})$
8       $e_{j}=\arg\max_{e\in A_{j}}\hat{F}^{j}_{S_{j}}(e)$
9       $S_{j+1}=S_{j}+e_{j}$
10
11 end for
return $S_{k+1}$
Algorithm 4 Mini-batch greedy
In Section 2 we analyze the relation between the batch parameter, $\alpha$, and the the type of approximate incremental oracles guaranteed by our sampling procedure. We state the main theorem for the section below.
Theorem 3.
The mini-batch greedy algorithm (Algorithm 4) maximizing a non-negative monotone submodular function has the following guarantees:
  • If $F$ has curvature bounded by $c$, and $\alpha=\Theta(\frac{\log n}{{\epsilon}^{2}(1-c)})$ it holds w.h.p that $\forall j\in[k]$ that $\hat{F}^{j}_{S_{j}}$ is an $(1-\epsilon)$-approximate incremental oracle.
  • If $\alpha=\Theta({\epsilon}^{-2}\gamma\log n)$ it holds w.h.p that $\forall j\in[k]$ that $\hat{F}^{j}_{S_{j}}$ is an additive $(\epsilon/\gamma)$-approximate incremental oracle, for any parameter $\gamma>0$.
Furthermore, the number of oracle evaluations during preprocessing is $O(nN)$ and an expected $\alpha(\sum_{i=1}^{N}p_{i})(\sum_{j=1}^{k}\left|A_{j}\right|)=O(\alpha kn^{2})$ during execution.
Combining Theorem 3 (setting $\gamma=k$ for a cardinality constraint and $\gamma=kp$ for a $p$-system constraint) with Theorem 1 and Theorem 2 we state our main result.
Theorem 4.
The mini-batch greedy algorithm maximizing a non-negative monotone submodular function requires $O(nN)$ oracle calls during preprocessing and has the following guarantees:
  • If $F$ has curvature bounded by $c$, it achieves w.h.p a $(1-1/e-{\epsilon})$-approximation under a cardinality constraint and $(\frac{1-{\epsilon}}{1+p})$-approximation under a $p$-system constraint with an expected $O(\frac{kn^{2}\log n}{{\epsilon}^{2}(1-c)})$ oracle evaluations for both cases.
  • It achieves w.h.p a $(1-1/e-{\epsilon})$-approximation under a cardinality constraint and $(\frac{1-{\epsilon}}{1+p})$-approximation under a $p$-system constraint with an expected $O(k^{2}(n/{\epsilon})^{2}\log n)$ and $O(k^{2}p(n/{\epsilon})^{2}\log n)$ oracle evaluations respectively.
1.2 Related work
Approximate oracles
Apart from the results of (Goundan and Schulz, 2007; Călinescu et al., 2011) there are works that use different notions of an approximate oracle. Several works consider an approximate oracle $\hat{F}$, such that $\forall S\subseteq E,\left|\hat{F}(S)-F(S)\right|<{\epsilon}F(S)$ (Crawford et al., 2019; Horel and Singer, 2016; Qian et al., 2017). The main difference of these models to our work is the fact that they do not assume the surrogate function, $\hat{F}$, to be submodular. This adds a significant complication to the analysis and degrades the performance guarantees.
Mini-batch methods
The closest results resembling mini-batch methods for submodular functions are due to Buchbinder et al. (2015); Mirzasoleiman et al. (2015). They improve the expected query complexity of the greedy algorithm under a cardinality constraint by only considering a small random subset of $E\setminus S_{j}$ at the $j$-th iteration. We note that their approach can be combined into our mini-batch algorithm, reducing our query complexity by a $\tilde{\Theta}(k)$ factor, resulting in an approximation guarantee in expectation instead of w.h.p.
2 Analysis of the mini-batch greedy algorithm
Let us start by bounding the expected size of $\hat{F}$ in Algorithm 3. We start with the following useful lemma proved in Kenneth and Krauthgamer (2023). We restate the proof for completeness.
Lemma 5.
The expected size of $\hat{F}$ is $\alpha\sum_{i=1}^{N}{p_{i}}\leq\alpha n$.
Proof.
We start by showing that $\sum_{i=1}^{N}p_{i}\leq n$. Let us divide the range $[N]$ into
$A_{e}=\left\{i\in N\mid e=\operatorname*{arg\,max}_{e^{\prime}\in E,F_{\emptyset}(e^{\prime})\neq 0}\frac{f^{i}_{\emptyset}(e^{\prime})}{F_{\emptyset}(e^{\prime})}\right\}$
If 2 elements in $E$ achieve the maximum value for some $i$, we assign it to a single $A_{e}$ arbitrarily.
$\sum_{i=1}^{N}p_{i}=\sum_{i=1}^{N}\max_{e\in E}\frac{f^{i}_{\emptyset}(e)}{F_{\emptyset}(e)}=\sum_{e\in E}\sum_{i\in A_{e}}\frac{f^{i}_{\emptyset}(e)}{F_{\emptyset}(e)}=\sum_{e\in E}\frac{\sum_{i\in A_{e}}f^{i}_{\emptyset}(e)}{F_{\emptyset}(e)}\leq\sum_{e\in E}1\leq n$
Let $X_{i}$ be an indicator variable for the event $w_{i}>0$. We are interested in $\sum_{i=1}^{N}\mathbb{E}[X_{i}]$. It holds that:
$\sum_{i=1}^{N}\mathbb{E}[X_{i}]=\sum_{i=1}^{N}\alpha_{i}\leq\sum_{i=1}^{N}\alpha p_{i}=\alpha\sum_{i=1}^{N}p_{i}\leq\alpha n$
Next, let us show that $\hat{F}$ returned by Algorithm 3 is indeed an (additive) approximate incremental oracle w.h.p. We make use of the following Chernoff bound.
Theorem 6 (Chernoff bound (Motwani and Raghavan, 1995)).
Let $X_{1},...,X_{N}$ be independent random variables in the range $[0,a]$. Let $X=\sum_{i=1}^{N}X_{i}$. Then for any ${\epsilon}\in[0,1]$ and $\mu\geq\mathbb{E}[T]$,
$\mathbb{P}(|X-\mathbb{E}[X]|\geq{\epsilon}\mu)\leq 2\exp\left(-\frac{{\epsilon}^{2}\mu}{3a}\right)$
The following lemma provides concentration guarantees for $\hat{F}$ in Algorithm 3.
Lemma 7.
For every $S\subseteq E$ ($\hat{F}$ sampled after $S$ is fixed) and for every $e\in E$ and $\mu\geq F_{S}(e)$, it holds that
$\mathbb{P}\left[|\hat{F}_{S}(e)-F_{S}(e)|\geq{\epsilon}\mu\right]\leq 2\exp\left(-\frac{{\epsilon}^{2}\mu}{3F_{\emptyset}(e)/\alpha}\right)$
Proof.
Fix some $e\in E$. Let $G=\sum_{i\in I}f^{i}$, where $I=\{i\in[N]\mid\alpha_{i}=1\}$. Let $F^{\prime}_{S}(e)=F_{S}(e)-G_{S}(e)$ and $\hat{F}^{\prime}_{S}(e)=\hat{F}_{S}(e)-G_{S}(e)$. Let $J=[N]\setminus I$. It holds that:
$\mathbb{P}\left[|\hat{F}_{S}(e)-F_{S}(e)|\geq{\epsilon}\mu\right]=\mathbb{P}\left[|\hat{F}^{\prime}_{S}(e)+G_{S}(e)-F^{\prime}_{S}(e)-G_{S}(e)|\geq{\epsilon}\mu\right]=\mathbb{P}\left[|\hat{F}^{\prime}_{S}(e)-F^{\prime}_{S}(e)|\geq{\epsilon}\mu\right]$
Due to the fact that $\mathbb{E}\left[{w_{i}}\right]=1$ we have $\mathbb{E}[\hat{F}^{\prime}_{S}(e)]=\mathbb{E}[\sum_{i\in J}w_{i}f^{i}_{S}(e)]=F^{\prime}_{S}(e)$. As $f^{i}$’s are monotone, it holds that $\mu\geq F_{S}(e)\geq F^{\prime}_{S}(e)$. Applying a Chernoff bound (Theorem 6) we have
$\mathbb{P}\left[|\hat{F}^{\prime}_{S}(e)-F^{\prime}_{S}(e)|\geq{\epsilon}\mu\right]\leq 2\exp\left(-{\epsilon}^{2}\mu/3a\right)$
where $a=\max\{w_{i}f^{i}_{S}(e)\}_{i\in J}$. Recall that $w_{i}=1/\alpha_{i}$ where $\alpha_{i}=\min\{1,\alpha p_{i}$} and $\alpha_{i}<1$ for all $i\in J$. Let us upper bound $a$.
$a$ $=\max_{i\in J}w_{i}f^{i}_{S}(e)=\max_{i\in J}\frac{f^{i}_{S}(e)}{\alpha p_{i}}=\max_{i\in J}\frac{f^{i}_{S}(e)}{\alpha\cdot\underset{e^{\prime}\in E}{\max}\frac{f^{i}_{\emptyset}(e^{\prime})}{F_{\emptyset}(e^{\prime})}}\leq\max_{i\in J}\frac{f^{i}_{\emptyset}(e)}{\alpha\cdot\frac{f^{i}_{\emptyset}(e)}{F_{\emptyset}(e)}}=\frac{F_{\emptyset}(e)}{\alpha}$
Where the inequality is due to submodularity and non-negativity in the numerator and maximality in the denominator. Given the above upper bound for $a$ we get:
$\mathbb{P}\left[|\hat{F}_{S}(e)-F_{S}(e)|\geq{\epsilon}\mu\right]\leq 2\exp\left(-{\epsilon}^{2}\mu/3a\right)\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha\mu}{3F_{\emptyset}(e)}\right)$
Using the above we state the main theorem for this section. See 3
Proof.
The number of oracle evaluations is due to Lemma 5 and the fact that the algorithm executes for $k$ iteration and must evaluate $\left|A_{j}\right|\leq n$ elements per iteration.
Let us prove the approximation guarantees. Let us start with the bounded curvature case. Fix some $S_{j}$. As $\hat{F^{j}}$ is sampled after $S_{j}$ is fixed, we can fix some $e\in E$ and apply Lemma 7 with $\mu=F_{S_{j}}(e)$. We get that:
$\mathbb{P}\left[|\hat{F}^{j}_{S_{j}}(e)-F_{S_{j}}(e)|\geq{\epsilon}F_{S_{j}}(e)\right]\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha F_{S_{j}}(e)}{3F_{\emptyset}(e)}\right)\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha(1-c)}{3}\right)\leq 1/n^{3}$
Where the second inequality is due to the fact that $F_{S_{j}}(e)/F_{\emptyset}(e)\geq\min_{S\subseteq E,e^{\prime}\in E\setminus S}F_{S}(e^{\prime})/F_{\emptyset}(e^{\prime})=1-c$, and the last transition is by setting an appropriate constant in $\alpha=\Theta(\frac{\log(n)}{{\epsilon}^{2}(1-c)})$.
When the curvature is not bounded, we break the analysis into cases.
$F_{\emptyset}(e)\leq\gamma F_{S_{j}}(e)$:
Setting $\mu=F_{S_{j}}(e)$ we get:
$\mathbb{P}\left[|\hat{F}^{j}_{S_{j}}(e)-F_{S_{j}}(e)|\geq{\epsilon}F_{S_{j}}(e)\right]\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha F_{S_{j}}(e)}{3F_{\emptyset}(e)}\right)\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha}{3\gamma}\right)\leq 1/n^{3}$
$F_{\emptyset}(e)>\gamma F_{S_{j}}(e)$:
Here we can set $\mu=F_{\emptyset}(e)/\gamma\geq F_{S_{j}}(e)$ and get:
$\mathbb{P}\left[|\hat{F}^{j}_{S_{j}}(e)-F_{S_{j}}(e)|\geq{\epsilon}F_{\emptyset}(e)/\gamma\right]\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha F_{\emptyset}(e)}{3\gamma F_{\emptyset}(e)}\right)\leq 2\exp\left(-\frac{{\epsilon}^{2}\alpha}{3\gamma}\right)\leq 1/n^{3}$
Where in both cases the last inequality is by setting an appropriate constant in $\alpha=\Theta(\epsilon^{-2}\gamma\log n)$. Note that in the first case we get an $(1-\epsilon)$-approximate incremental oracle and in the second an additive $({\epsilon}/\gamma)$-approximate incremental oracle. The second case is the worse of the two for our analysis.
For both bounded and unbounded curvature, we take a union bound over all $e\in E$ and $j\in[k]$ (at most $n^{2}$ values), which concludes the proof. ∎
Note that in the above we use the fact that $F_{\emptyset}(e)\leq F(S^{*})$ (recall that $S^{*}$ is the optimal solution for $F$) to get the second result. This is sufficient for our proofs to go through, however, the theorem has a much stronger guarantee which might be useful in other contexts.
3 Experiments
We perform experiments on the following datasets:
Uber pickup
This dataset consists of Uber pickups in New York city in May 2014555 https://www.kaggle.com/fivethirtyeight/uber-pickups-in-new-york-city. The set contains $652,434$ records, where each record consists of a longitude and latitude, representing a pickup location. Following (Rafiey and Yoshida, 2022) we aim to find $k$ positions for idle drivers to wait from a subset of popular pickup locations. We formalize the problem as follows. We run Lloyd’s algorithm on the dataset, $X$, and find 100 cluster centers. We set these cluster centers to be the ground set $E$. We define the goal function $F:2^{E}\rightarrow\mathbb{R}^{+}$ as $F(S)=\sum_{v\in X}f_{v}(S)$ where $f_{v}(S)=\max_{e\in E}d(v,e)-\min_{e\in S}d(v,e)$, where $d(v,e)$ is the Manhattan distance between $v$ and $e$. The goal of the problem is to maximize $F(S)$ subject to $\left|S\right|\leq k$.
Discogs (Kunegis, 2013)
This dataset [^6] http://konect.cc/networks/discogs_lstyle/ provides audio record information structured as a bipartite graph $G=(L,R;E^{\prime})$. The left nodes represent labels and the right nodes represent styles. Each edge $(u,v)\in L\times R$ signifies the involvement of a label $u$ in producing a release of a style $v$. The dataset comprises $|R|=383$ labels, $|L|=243,764$ styles, and $|E^{\prime}|=5,255,950$ edges. We aim to select $k$ styles that cover the activity of the maximum number of labels. For problem formalization, a maximum coverage function $F:2^{R}\rightarrow\mathbb{R}$ is constructed, where $F(S)=\sum_{v\in L}f_{v}(S)$, and $f_{v}(S)$ equals 1 if $v$ is adjacent to some element of $S$ and 0 otherwise. The objective is to maximize $F(S)$ subject to $|S|\leq k$.
Examplar-based clustering
We consider the problem of selecting representative subset of $k$ images from a massive data set. We present experiments for both the CIFAR100 and the FashionMNIST datasets. For each dataset we consider a subset of $50,000$ images, denoted by $X$. We flatten every image into a one dimensional vector, subtract from it the mean of all images and normalize it to unit norm. We take the distance between two elements in $X$ as $d(x,x^{\prime})=\|x-x^{\prime}\|^{2}$. Here the ground set is simply the dataset, $E=X$. Similarly to the Uber pickup dataset we define the goal function $F:2^{E}\rightarrow\mathbb{R}^{+}$ as $F(S)=\sum_{v\in X}f_{v}(S)$ where $f_{v}(S)=\max_{e\in E}d(v,e)-\min_{e\in S}d(v,e)$. The goal of the problem is to maximize $F(S)$ subject to $\left|S\right|\leq k$.
Experimental setup
We compare the sparsifier approach with the mini-batch approach for each of the above datasets as follows. We first compute the $p_{i}$’s and fix a parameter $\beta\in(0,1)$. We take $\alpha$ in the algorithm such that that $\alpha\sum_{i}^{N}p_{i}=\beta N$. That is, $\beta$ is the desired fraction of the dataset we wish to sample for the sparsifier / mini-batch. Note that the $p_{i}$’s are the same for the sparsifier and the minibatch approach, and the only difference is whether we sample one time in the beginning (sparsifier) or for every iteration of the greedy algorithm (mini-batch). Parameterizing by $\beta$ guarantees that both algorithms make the same number of oracle calls in expectation, and it is sufficient to only compare the performance. For example, setting $\beta=0.01$ means that the sparsifier will sample 1% of the data and execute the greedy algorithm, while the mini-batch algorithm will sample 1% of the data for every iteration of the greedy algorithm. This also means that the number of queries made for $\beta=0.01$ is 1% of the number of queries made by the full-batch algorithm.
We present our results in Figure 1 where we plot results for different $\beta$ values for both the sparsifier and the mini-batch approach for different values of $k$. Every point on the graph is the average of 20 executions with the same $k,\beta$. The $y$-axis is the relative utility of the solution compared to the (full-batch) greedy algorithm. It is clear from our experiments that the mini-batch approach is superior to the sparsifier approach for small values of $\beta$.
(a) Uber pickup
(b) Discogs
(c) Cifar100
(d) FashionMNIST
Figure 1: Comparative visualizations
Appendix A Proof of Theorem 2
Theorem 2 directly follows from the two lemmas below.
Lemma 8.
Let ${\epsilon}^{\prime}\leq{\epsilon}/2k$. Algorithm 2 with an additive ϵsuperscriptitalic-${\epsilon}^{\prime}$-approximate incremental oracle achieves a $(1-1/e-{\epsilon})$-approximation under a cardinality constraint $k$.
Proof.
Let $S^{*}$ be some optimal solution for $F$. We start by proving that the following holds for every $j\in[k]$:
$F(S_{j+1})-F_{(}S_{j})\geq\frac{1}{k}((1-{\epsilon})F(S^{*})-F(S_{j}))$
Fix some $j\in[k]$ and let $S^{*}\setminus S_{j}=\left\{e^{*}_{1},\dots,e^{*}_{\ell}\right\}$ where $\ell\leq k$. Let $S^{*}_{t}=\left\{e^{*}_{1},\dots,e^{*}_{t}\right\}$, and $S^{*}_{0}=\emptyset$. Let us first use submodularity and monotonicity to upper bound $F(S^{*})$.
$F(S^{*})\leq F(S^{*}+S_{j})=F(S_{j})+\sum_{t=1}^{\ell}[F(S_{j}+S^{*}_{t})-F(S_{j}+S^{*}_{t-1})]$
$\leq F(S_{j})+\sum_{t=1}^{\ell}F_{S_{j}}(e^{*}_{t})\leq F(S_{j})+\sum_{t=1}^{\ell}\max_{e\in E\setminus S_{j}}F_{S_{j}}(e)$
$\leq F(S_{j})+k\max_{e\in E\setminus S_{j}}F_{S_{j}}(e)\leq F(S_{j})+k(\max_{e\in E\setminus S_{j}}\hat{F}^{j}_{S_{j}}(e)+{\epsilon}^{\prime}F(S^{*}))$
Where the last inequality is due to the fact that $\hat{F}^{j}_{S_{j}}$ is an additive ${\epsilon}^{\prime}$-approximate incremental oracle.
Noting that $e_{j}=\operatorname*{arg\,max}_{e\in E\setminus S_{j}}\hat{F}^{j}_{S_{j}}(e)$ we get that:
$F(S^{*})\leq F(S_{j})+k(\hat{F}^{j}_{S_{j}}(e_{j})+{\epsilon}^{\prime}F(S^{*}))$
$\implies\hat{F}^{j}_{S_{j}}(e_{j})\geq\frac{1}{k}((1-{\epsilon}^{\prime}k)F(S^{*})-F(S_{j}))$
The above lower bounds the progress on the $j$-th mini-batch. Now, let us bound the progress on $F$. Again, we use the fact that $\hat{F}^{j}_{S_{j}}$ is an additive ${\epsilon}^{\prime}$-approximate incremental oracle.
$F(S_{j+1})-F(S_{j})\geq\hat{F}^{j}_{S_{j}}(e_{j})-{\epsilon}^{\prime}F(S^{*})$
$\geq\frac{1}{k}((1-{\epsilon}^{\prime}k)F(S^{*})-F(S_{j}))-{\epsilon}^{\prime}F(S^{*})\geq\frac{1}{k}((1-2{\epsilon}^{\prime}k)F(S^{*})-F(S_{j}))$
Finally, using the fact that ${\epsilon}^{\prime}\leq{\epsilon}/2k$ we get:
$F(S_{j+1})-F(S_{j})\geq\frac{1}{k}((1-{\epsilon})F(S^{*})-F(S_{j}))$
Rearranging, the result directly follows using standard arguments.
$F(S_{k+1})>\frac{(1-{\epsilon})}{k}F(S^{*})+(1-\frac{1}{k})F(S_{k})\geq\frac{(1-{\epsilon})}{k}F(S^{*})(\sum_{i=0}^{k}(1-\frac{1}{k})^{i})+F(\emptyset)$
$\geq F(S^{*})\frac{(1-{\epsilon})(1-\frac{1}{k})^{k}}{k(1-(1-\frac{1}{k}))}=(1-{\epsilon})(1-\frac{1}{k})^{k}F(S^{*})\geq(1-{\epsilon})(1-1/e)F(S^{*})\geq(1-1/e-{\epsilon})F(S^{*})$
Lemma 9.
Let ${\epsilon}^{\prime}\leq{\epsilon}/2kp$. Algorithm 2 with an additive ϵsuperscriptitalic-${\epsilon}^{\prime}$-approximate incremental oracle achieves a $(\frac{1-{\epsilon}}{1+p})$-approximation under a $p$-system constraint.
Proof.
Let $S^{*}$ be some optimal solution for $F$. Assume without loss of generality that the solution returned by the algorithm consists of $k$ elements $S_{k+1}=\left\{e_{1},\dots,e_{k}\right\}$.
We show the existence of a partition $S^{*}_{1},S^{*}_{2},\dots,S^{*}_{k}$ of $S^{*}$ such that $F_{S_{j}}(e_{j})\geq\frac{1}{p}F_{S_{k+1}}(S^{*}_{j})-2{\epsilon}^{\prime}F(S^{*})$. Note, we allow some of the sets in the partition to be empty.
Define $T_{k}=S^{*}$. For $j=k,k-1,...,2$ execute: Let $B_{j}=\left\{e\in T_{j}\mid S_{j}+e\in\mathcal{I}\right\}$. If $\left|B_{j}\right|\leq p$ set $S^{*}_{j}=B_{j}$; else pick an arbitrary $S^{*}_{j}\subset B_{j}$ with $\left|S^{*}_{j}\right|=p$. Then set $T_{j-1}=T_{j}\setminus S^{*}_{j}$ before decreasing $j$. After the loop set $S^{*}_{1}=T_{1}$. It is clear that for $j=2,...,k,\left|S^{*}_{j}\right|\leq p$.
We prove by induction over $j=0,1,...,k-1$ that $\left|T_{k-j}\right|\leq(k-j)p$. For $j=0$, when the greedy algorithm stops, $S_{k+1}$ is a maximal independent set contained in $E$, therefore any independent set (including $T_{k}=S^{*}$) satisfies $\left|T_{k}\right|\leq p\left|S_{k+1}\right|=pk$. We proceed to the inductive step for $j>0$. There are two cases: (1) $\left|B_{k-j+1}\right|>p$, which implies that $\left|S^{*}_{k-j+1}\right|=p$ and using the induction hypothesis we get that $\left|T_{k-j}\right|=\left|T_{k-j+1}\right|-\left|S^{*}_{k-j+1}\right|\leq(k-j+1)p-p=(k-j)p$. (2) $\left|B_{k-j+1}\right|\leq p$, it holds that $T_{k-j}=T_{k-j+1}\setminus B_{k-j+1}$. Let $Y=S_{k-j+1}+T_{k-j}$. Due to the definition of $B_{k-j+1}$ it holds that $S_{k-j+1}$ is a maximal independent set in $Y$. It holds that $T_{k-j}$ is independent and contained in $Y$, therefore $\left|T_{k-j}\right|\leq p\left|S_{k-j+1}\right|=p(k-j)$.
Finally, we get that $\left|T_{1}\right|=\left|S^{*}_{1}\right|\leq p$. By construction it holds that $\forall j\in[k],\forall e\in S^{*}_{j},S_{j}+e\text{ is independent}$. From the choice made by the greedy algorithm and the fact that $\hat{F}^{j}_{S_{j}}$ is an additive ${\epsilon}^{\prime}$-approximate incremental oracle it follows that for each $e\in S^{*}_{j}$:
$F_{S_{j}}(e_{j})\geq\hat{F}^{j}_{S_{j}}(e_{j})-{\epsilon}^{\prime}F(S^{*})\geq\hat{F}^{j}_{S_{j}}(e)-{\epsilon}^{\prime}F(S^{*})\geq F_{S_{j}}(e)-2{\epsilon}^{\prime}F(S^{*})$
Hence,
$\left|S^{*}_{j}\right|F_{S_{j}}(e_{j})\geq\sum_{e\in S^{*}_{j}}(F_{S_{j}}(e)-2{\epsilon}^{\prime}F(S^{*}))\geq F_{S_{j}}(S^{*}_{j})-2{\epsilon}^{\prime}\left|S^{*}_{j}\right|F(S^{*})\geq F_{S_{k+1}}(S^{*}_{j})-2{\epsilon}^{\prime}\left|S^{*}_{j}\right|F(S^{*})$
Using submodularity in the last two inequalities.
For all $j\in\left\{1,2,...,k\right\}$ it holds that $\left|S^{*}_{j}\right|\leq p$, and thus $F_{S_{j}}(e_{j})\geq\frac{1}{p}F_{S_{k+1}}(S^{*}_{j})-2{\epsilon}^{\prime}F(S^{*})$. Using the partition we get that:
$F(S_{k+1})\geq\sum_{j=1}^{k}F_{S_{j}}(e_{j})\geq\sum_{j=1}^{k}(\frac{1}{p}F_{S_{k+1}}(S^{*}_{j})-2{\epsilon}^{\prime}F(S^{*}))$
$\geq\frac{1}{p}F_{S_{k+1}}(S^{*})-2{\epsilon}^{\prime}kF(S^{*})\geq\frac{1}{p}(F(S^{*})-F(S_{k+1}))-2{\epsilon}^{\prime}kF(S^{*})$
Where the second to last inequality is due to submodularity and the last is due to monotonicity. Rearranging we get that:
$F(S_{k+1})\geq\frac{(1-2p{\epsilon}^{\prime}k)}{p+1}F(S^{*})$
As ${\epsilon}^{\prime}<\frac{{\epsilon}}{2pk}$ we get the desired result. ∎